# Skrull: Towards Efficient Long Context Fine-tuning through Dynamic Data Scheduling

### v1

@@ 1 Introduction
Long-context capabilities are important for large language models (LLMs) to handle various tasks such as long document summarization, question answering, multi-turn dialogue and code generation. Mainstream LLMs such as Llama [21, 10], Qwen [19] and GPT-4 [18] can support the context window of up to 128K tokens. Google’s Gemini [9] can even achieve up to 1M tokens per context window. Typically, additional training phases like long-context supervised fine-tuning (Long-SFT) as well as long-context continue pre-training (Long-CPT) are employed to extend the context length. For example, Llama3 [10] is fine-tuned with 99.89% short sequence (averaging under 1K tokens) and 0.11% long sequence (averaging around 37K tokens). Qwen2.5-Turbo [19] gradually extends context length by training on 40% long sequences and 60% short sequences. Training on those meticulous gathered datasets enables smoothly adaptation of LLMs to longer context while still maintaining the performance on short context tasks.
However, this heterogeneous data distribution in Long-SFT poses significant challenges for existing distributed LLM training frameworks [14, 20, 15], exhibiting sub-optimal efficiency. For instance, the heterogeneous data distribution poses a dilemma for parallelism and memory-reduction strategies. Specifically, long sequences necessitate context parallelism and other memory-reduction approaches due to their tremendous memory requirements. However, those approaches compromise the training efficiency for short ones due to the overheads like unnecessary communication and GPU under-utilization. Moreover, the wide sequence length distribution in long-SFT worsen the mismatch of computation characteristics in Attention module, which exhibit quadratic computational complexity and linear memory consumption [7, 6], leading to another dilemma for load balance problem.
@@ 2 Preliminaries
Data parallelism [15, 25, 20] partitions the training samples to multiple workers and each worker maintains a complete model weight replica. In each iteration, all workers process a subset of global batch independently and then synchronize the gradients across all DP ranks. However, due to the inherent synchronization semantic in DP, the load balance becomes a noticeable problem, especially in long context scenarios.
@@
Context parallelism partitions the input tensor along the sequence length dimension and distributes it to multiple workers [12, 16, 10]. CP is emerging as an inevitable parallel strategy when handling long context. In the Transformer architecture, the primary challenge of CP stems from the parallelization of Attention module because each tokens needs to attend to other tokens in the sequence. Consequently, the communication in CP is inevitable. Notably, DACP, proposed in this paper, leverages data scheduling to minimize the overheads caused by CP and is orthogonal to specific CP implementations.
@@ 3.1 Heterogeneous Sequence Length Distribution
As shown in Figure 1(a), we observe pronounced variance in the sequence length distribution across real-world Long-SFT datasets, including Wikipedia [2], LMsysChat1M [26] and ChatQA2-Long-SFT [1]. Among them, the sequence length distribution of ChatQA2-long-SFT exhibits a bimodal pattern, where the proportions of long and short sequences are nearly equal. Specifically, approximately 40% of sequences are shorter than 8K tokens, while the remaining 60% exceed this threshold. As comparison, long-tail distributions represent another typical pattern in Long-SFT datasets. In Llama3’s internally collected Long-SFT datasets [10], we find that 99.89% of sequences are under 1K tokens on average, while the remaining 0.11% are approximately 37K tokens, showcasing extremely skewed long-tail distribution. Due to data accessibility constraints, we plot the sequence length distribution of Wikipedia and LMsysChat1M in Figure 1(a), which have the identical feature with Llama3’s Long-SFT dataset. Table 1 lists the portions under different lengths thresholds for these three datasets, highlighting the heterogeneous sequence length distribution in Long-SFT.
@@ 3.2 Performance Degradations for Short Sequences
In this section, we discuss our observation on the performance degradations and GPU under-utilization for short sequences in Long-SFT training. During the training process, the context parallelism degree and other memory reduction strategies such as gradient accumulation are set to accommodate the longest sequence in datasets to avoid out-of-memory errors (OOMs). However, these training settings degrade their performance for the shorter sequences, which make up the majority in Long-SFT datasets. As shown in Figure 1(b), we test the performance of Attention module [6] under different CP degrees. Results demonstrate, especially for the short sequences, higher CP degree exacerbates kernel execution efficiency. Additionally, context parallelism also brings unnecessary communication overhead to short sequences. Also, the memory reduction strategies tailored to long sequence lead to low GPU memory utilization for the most time.
@@ 4.1 Distributed-aware Context Parallelism
To simultaneously achieve high efficiency for all the sequences, we propose distributed-aware context parallelism (DACP). As shown in Figure 2(c), DACP dynamically determines whether to distribute the sequences to avoid unnecessary overheads or not. On the one hand, DACP preserves the original context parallel settings to maintain the ability of handling long sequences. On the other hand, DACP selectively schedules short sequences entirely within a single device to minimize the degradation. Therefore, based on distinct computational characteristics, DACP classifies sequences into two categories: (i) distributed sequences requiring context parallelism, and (ii) local sequences needing efficient processing and intended to reside entirely within a single device. Notably, these sequences are still processed within a shared CP group without increasing the number of GPUs used in training. Furthermore, as illustrated in Figure 2(d), DACP brings an additional opportunity to overlap the communication of distributed sequences and the computation of local sequences in Attention module due to the inherent independence between distributed and local sequences.
However, the scheduling process presents significant challenges. First, inappropriate sequence classification may lead to out-of-memory errors (OOMs). Second, the local sequences are varying in length and pose load imbalance issue across CP ranks. To fully explore the relationship between scheduling plans and performance gains, we first analyze the computation and memory features in Appendix A. Through offline profiling, we model the computation (see in Appendix A.2) by $FLOPs$ function and latency estimation function $T_{comp}$. Additionally, we map the sequence length to the memory consumption and derive the BucketSize $C$ which indicates the capacity of total sequence length per ranks. The BucketSize $C$ plays a vital role in measuring the memory constrain during Skrull’s scheduling. More details are listed in Appendix A.1. Similarly, we model communication volume function $Volume$ and latency function $T_{comm}$, as detailed in Appendix A.3. Finally, we formulate the scheduling process as an optimization problem as follows. The frequently used notions are listed in Table 2.
@@
$N$ CP degrees. $FLOPs$ FLOPs estimation function.
@@
$B_{k}ij$ Assign $k$-th sequence to $i-th$ DP rank and $j-th$ micro-batch. $ws$ DP degree
$Volume$ Communication volume count function.
@@
We first define the sequence classification array $D\in\{0,1\}^{K}$ ( 0 for local sequence and 1 for distributed sequence) and local sequence assignment matrix $P\in\{0,1\}^{K\times N}$ ( 1 for assignment and 0 for not). For example, $D_{k}=1$ indicates that the $k-th$ sequence with length of $S_{k}$ is scheduled to be computed in a distributed manner. Similarly, $P_{kj}=1$ indicates that the $k-th$ sequence is assigned to device $j$, implying $D_{k}=0$. Given a micro-batch comprising $K$ sequences with lengths $S_{k}$ ($k=(0,\ldots,K-1)$), BucketSize $C$ and CP degree $N$, the scheduling process of DACP can be formulated as follows:
@@
$\text{Time}_{j}=\max\left(T_{comm}\left(\text{V}\right),T_{comp}\left(\text{Local}_{j}\right)\right)+T_{comp}\left(\text{Dist}\right),\quad\forall j$ (2)
$\text{Local}_{j}=\sum_{k}\text{FLOPs}\left(P_{kj}\cdot S_{k}\right),\quad\forall j$ (3)
@@
$\sum_{j}P_{kj}+D_{k}=1,\quad\forall k$ (6)
$\sum_{k}S_{k}\cdot P_{kj}+\frac{D_{k}\cdot S_{k}}{\text{N}}\leq C,\quad\forall j$ (7)
Here, our optimization goal is to find the optimal $D$ and $P$ to minimize $TDACP$, which indicates the total duration in one micro-batch. As shown in Equation 1, $TDACP$ is determined by the maximum execution time $\text{Time}_{j}$ across all CP ranks $j$. Specifically, as described in Equation 2, $\text{Time}_{j}$ consists of two components: (1) the overlapping term, defined as the maximum of the communication time $T_{comm}(\text{V})$ and the computation time $T_{comp}(\text{Local}_{j})$ for local sequences, and (2) the computation time $T_{comp}(\text{Dist})$ for distributed sequences. Here, $T_{comm}$ depends on the communication volume V, as modeled in Equation 5. Similarly, $T_{comp}$ utilizes the results from Equations 3 and 4, which compute the FLOPs for local sequences on CP rank $j$ and distributed sequences, respectively. Finally, Equation 6 ensures the completeness of data scheduling, while Equation 7 enforces the memory constraint.
@@ 4.2 Global Data Scheduling
First, the heterogeneous sequence length distribution also leads the to severe load imbalance across different micro-batches, resulting in the sub-optimal training efficiency in Long-SFT scenarios. Second, to achieve maximum performance gains in DACP, meticulous micro-batching strategy is essential. For example, pairing long and short sequences with appropriate memory pressure can expand the valid scheduling space for DACP. Specifically, micro-batches with large total sequence lengths increase the risk of OOMs and limit the optimizations in DACP, such as selective sharding. In contrast, micro-batches with small total sequence lengths introduce GPU under-utilization, degrading the end-to-end performance. Therefore, as shown in Figure 2(b), Skrull employs Global Data Scheduling (GDS), which derives the optimal micro-batching strategy from the global batch. We limit the scheduling scope to the global batch because it represents the maximum scope that maintains mathematical equivalence for mainstream optimizers such as Adam [13] and AdamW [17].
@@
$\sum_{ij}B_{kij}=1,\quad\forall k$ (9)
$\sum_{ki}B_{kij}*S_{k}\leq C*N,\quad\forall j$ (10)
$\text{Time}_{ij}=\textbf{TDACP}\left(B_{kij}*S_{k}\right),\quad\forall k$ (11)
@@ 4.3 Efficient Online Scheduling
Although some solvers like [4] can derive the the optimal scheduling plan, its long solving time makes it impractical for scheduling during runtime. To achieve online scheduling during Long-SFT, we resort to design lightweight heuristic scheduling algorithm. Notably, our scheduling algorithm is integrated into the DataLoader and introduces near-zero overhead to the training process.
@@
Unfortunately, we cannot balance the computation and memory at the same time. The reason is that, after applying FlashAttention [7, 6], the correlation between computation complexity and and sequence length ($n$) is $O(n^{2})$, however, the correlation between memory is $O(n)$. Moreover, with the sequence length increasing, the portion of Attention module gradually dominates the computation load, making it more difficult to balance the computation and memory. Worse still, the model configuration (e.g., KV heads, hidden size) also impacts. Due to the limited page, we list the details in Appendix A.
@@ 4.3.2 Heuristics
We first summarize three principles of algorithm design in DACP. (i) Avoid sharding: We strive to avoid sequence sharding and assume that all sequences will be handled locally first. (ii) Prioritize computation: We prioritize balancing computation over memory to achieve better performance. (iii) Roll-back mechanism: We continuously monitor the estimated memory consumption and revert decisions when necessary. The roll-back mechanism guarantees the memory constrains outlined in Equation 7 and Equation 10, while enabling more aggressive scheduling attempts based on (i) and (ii). Our heuristic for DACP is listed in Algorithm 1. Given a micro-batch containing $K$ sequences with lengths $S[K]$ and a predefined BucketSize $C$, the algorithm outputs the sequence classification and assignment results in the form of an array $ret$. In this array, a value of -1 at the i-th position indicates that the i-th sequence is to be sharded, while a value $v=(0,\ldots,ws-1)$ indicates that the i-th sequence is assigned to CP rank $v$ entirely. To better balance computation while ensuring memory constrains, we maintain two arrays during DACP scheduling: RemainBucket $RB$ and Loads $L$, which represent the current memory budget and computation load, respectively. We first sort the sequences in ascending order. For the each sequence, we sequentially assign it to the bucket (as well as CP rank) with minimum $L$ to avoid sharding and prioritize balancing computation (line 6-8). If the bucket cannot accommodate the sequence, we attempt to assign it to the bucket with the maximum $RB$ to avoid sharding (line 10-12). If both attempts fail, we classify the sequence as a distributed sequence and attempt to shard it (line 14-16). However, if the bucket with minimum of $RB$ cannot handle the sub-sequence after sharding, this indicates that the earlier process incorrectly classified inappropriate sequences as local sequences within this bucket. To address this, we employ a roll-back mechanism (line 18 and Appendix B). This mechanism identifies a local sequence in the bucket, shards it to reduce memory pressure, and resumes the assignment process. If the roll-back fails due to the absence of local sequences in the bucket, we return a DACP scheduling error. In such cases, GDS will also revert the batching plan (see the Section 4.3.2). Notably, every assignment updates $RB$ and $L$ through the predefined functions UpdateLocal and UpdateAll. The details of these functions including RollBack are further elaborated in Appendix B.
@@
1:SeqNum $K$, SeqLens $S[K]$, BucketSize $C$, CP degree $N$
@@
4:for $i=0$ to $N-1$ do
5:    $RB[i]{\leftarrow}C,L[i]{\leftarrow}0$triangleright Initialization
@@
7:for $i=0$ to $K-1$ do
8:    $t\leftarrow\text{argmin}(L)$triangleright Find rank t with minimum workload
@@
18:               $ret[i]{\leftarrow}{-}1$, UpdateAll$(i)$triangleright Distribute the sequence
@@
20:               Assert RollBack$(t,RB,L)$
21:               $i{\leftarrow}i{-}1$triangleright Roll-back to avoid OOMs
@@
Algorithm 2 demonstrates the heuristic scheduling algorithm of GDS. Given a global batch containing $K$ sequences with lengths $S[K]$, DP world size $ws$ and DP rank $dp\_rank$, the algorithm returns the scheduling result $mbs$, which consists of multiple micro-batches as inputs for Algorithm 1. We summarize three principles in our algorithm design. (i) Prioritize computation: We prioritize balancing computation across DP workers. To achieve this, we estimate the FLOPs (Appendix A.2) and employ a bin-packing algorithm to balance computational workloads at a coarse granularity (line 1). (ii) Pair long and short sequences: We sort the sequences within each DP rank and batch them in an interleaved manner (line 7). This approach ensures that long sequences are assigned more evenly across micro-batches. Additionally, each micro-batch contains several short sequences, enhancing both task overlapping and load balancing. (iii) Improve memory utilization: We estimate the total memory requirements and try to improve the concurrency with less number of micro-batches. Thanks to the roll-back mechanism (line 8), this method maximizes memory utilization while not increase the risk of OOMs. As shown in line 5, we gradually increase the number of micro-batches if the scheduling fails and requires a roll-back.
@@
1:SeqNum $K$, SeqLens $S[K]$, BucketSize $C$, CP degree $N$, DP WorldSize $ws$, DP_Rank $dp\_rank$
2:Micro-batches $mbs$
3:$Bin[ws]\leftarrow\text{Binpack}(ws,\text{FLOPs}(S[K]))$triangleright Coarse-fined balance
4:$Subset\leftarrow Bin[dp\_rank]$, $init\leftarrow\lceil\text{Sum}(Subset)/C\times N\rceil-1$
5:Sort($Subset$, ascending=True)
@@
8:    for $j\leftarrow 0$ to $init$ do
9:         $mbs.\text{append}(Subset[j::init])$triangleright Pair long and short sequences

### v2

@@ 1 Introduction
Long-context capabilities are important for large language models (LLMs) to handle various tasks such as long document summarization, question answering, multi-turn dialogue and code generation. Mainstream LLMs such as Llama [llama2_tech_report_2023, llama3_tech_report_2024], Qwen [qwen25_tech_report_2024] and GPT-4 [gpt4_tech_report_2024] can support the context window of up to 128K tokens. Google’s Gemini [gemini15_tech_report_2024] can even achieve up to 1M tokens per context window. Typically, additional training phases like long-context supervised fine-tuning (Long-SFT) as well as long-context continue pre-training (Long-CPT) are employed to extend the context length. For example, Llama3 [llama3_tech_report_2024] is fine-tuned with 99.89% short sequence (averaging under 1K tokens) and 0.11% long sequence (averaging around 37K tokens). Qwen2.5-Turbo [qwen25_tech_report_2024] gradually extends context length by training on 40% long sequences and 60% short sequences. Training on those meticulous gathered datasets enables smoothly adaptation of LLMs to longer context while still maintaining the performance on short context tasks.
However, this heterogeneous data distribution in Long-SFT poses significant challenges for existing distributed LLM training frameworks [megatron_v3_sp_selective_recompute_2023, deepspeed_zero_sc20, pytorch_distributed], exhibiting sub-optimal efficiency. For instance, the heterogeneous data distribution poses a dilemma for parallelism and memory-reduction strategies. Specifically, long sequences necessitate context parallelism and other memory-reduction approaches due to their tremendous memory requirements. However, those approaches compromise the training efficiency for short ones due to the overheads like unnecessary communication and GPU under-utilization. Moreover, the wide sequence length distribution in long-SFT worsen the mismatch of computation characteristics in Attention module, which exhibit quadratic computational complexity and linear memory consumption [flashattention_v1_2022, flashattention_v2_2024], leading to another dilemma for load balance problem.
@@ 2 Preliminaries
Data parallelism [pytorch_distributed, pytorch_fsdp, deepspeed_zero_sc20] partitions the training samples to multiple workers and each worker maintains a complete model weight replica. In each iteration, all workers process a subset of global batch independently and then synchronize the gradients across all DP ranks. However, due to the inherent synchronization semantic in DP, the load balance becomes a noticeable problem, especially in long context scenarios.
@@
Context parallelism partitions the input tensor along the sequence length dimension and distributes it to multiple workers [deepspeed_ulysess_2024, ring_attention_2024, llama3_tech_report_2024]. CP is emerging as an inevitable parallel strategy when handling long context. In the Transformer architecture, the primary challenge of CP stems from the parallelization of Attention module because each tokens needs to attend to other tokens in the sequence. Consequently, the communication in CP is inevitable. Notably, DACP, proposed in this paper, leverages data scheduling to minimize the overheads caused by CP and is orthogonal to specific CP implementations.
@@ 3.1 Heterogeneous Sequence Length Distribution
As shown in Figure 1(a), we observe pronounced variance in the sequence length distribution across real-world Long-SFT datasets, including Wikipedia [dataset_wikipedia], LMsysChat1M [dataset_lmsyschat1m] and ChatQA2-Long-SFT [dataset_chatqa2-long-sft-data_2025]. Among them, the sequence length distribution of ChatQA2-long-SFT exhibits a bimodal pattern, where the proportions of long and short sequences are nearly equal. Specifically, approximately 40% of sequences are shorter than 8K tokens, while the remaining 60% exceed this threshold. As comparison, long-tail distributions represent another typical pattern in Long-SFT datasets. In Llama3’s internally collected Long-SFT datasets [llama3_tech_report_2024], we find that 99.89% of sequences are under 1K tokens on average, while the remaining 0.11% are approximately 37K tokens, showcasing extremely skewed long-tail distribution. Due to data accessibility constraints, we plot the sequence length distribution of Wikipedia and LMsysChat1M in Figure 1(a), which have the identical feature with Llama3’s Long-SFT dataset. Table 1 lists the portions under different lengths thresholds for these three datasets, highlighting the heterogeneous sequence length distribution in Long-SFT.
@@ 3.2 Performance Degradations for Short Sequences
In this section, we discuss our observation on the performance degradations and GPU under-utilization for short sequences in Long-SFT training. During the training process, the context parallelism degree and other memory reduction strategies such as gradient accumulation are set to accommodate the longest sequence in datasets to avoid out-of-memory errors (OOMs). However, these training settings degrade their performance for the shorter sequences, which make up the majority in Long-SFT datasets. As shown in Figure 1(b), we test the performance of Attention module [flashattention_v2_2024] under different CP degrees. Results demonstrate, especially for the short sequences, higher CP degree exacerbates kernel execution efficiency. Additionally, context parallelism also brings unnecessary communication overhead to short sequences. Also, the memory reduction strategies tailored to long sequence lead to low GPU memory utilization for the most time.
@@ 4.1 Distributed-aware Context Parallelism
To simultaneously achieve high efficiency for all the sequences, we propose distributed-aware context parallelism (DACP). As shown in Figure 2(c), DACP dynamically determines whether to distribute the sequences to avoid unnecessary overheads or not. On the one hand, DACP preserves the original context parallel settings to maintain the ability of handling long sequences. On the other hand, DACP selectively schedules short sequences entirely within a single device to minimize the degradation. Therefore, based on distinct computational characteristics, DACP classifies sequences into two categories: (i) distributed sequences requiring context parallelism, and (ii) local sequences needing efficient processing and intended to reside entirely within a single device. Notably, these sequences are still processed within a shared CP group without increasing the number of GPUs used for training. Furthermore, as illustrated in Figure 2(d), DACP brings an additional opportunity to overlap the communication of distributed sequences and the computation of local sequences in Attention module due to the inherent independence between distributed and local sequences.
However, the scheduling process presents significant challenges. First, inappropriate sequence classification may lead to out-of-memory errors (OOMs). Second, the local sequences are varying in length and pose load imbalance issue across CP ranks. To fully explore the relationship between scheduling plans and performance gains, we first analyze the computation and memory features in Appendix C. Through offline profiling, we model the computation (see in Appendix C.2) by $FLOPs$ function and latency estimation function $T_{comp}$. Additionally, we map the sequence length to the memory consumption and derive the BucketSize $C$ which indicates the capacity of total sequence length per ranks. The BucketSize $C$ plays a vital role in measuring the memory constrain during Skrull’s scheduling. More details are listed in Appendix C.1. Similarly, we model communication volume function $Volume$ and latency function $T_{comm}$, as detailed in Appendix C.3. Finally, we formulate the scheduling process as an optimization problem as follows. The frequently used notions are listed in Table 2.
@@
NN CP degrees. FLOPsFLOPs FLOPs estimation function.
@@
$B_{k}ij$ Assign $k$-th sequence to ithi-th DP rank and jthj-th micro-batch. wsws DP degree
VolumeVolume Communication volume count function.
@@
We first define the sequence classification array $D\in\{0,1\}^{K}$ ( 0 for local sequence and 1 for distributed sequence) and local sequence assignment matrix $P\in\{0,1\}^{K\times N}$ ( 1 for assignment and 0 for not). For example, $D_{k}=1$ indicates that the kthk-th sequence with length of $S_{k}$ is scheduled to be computed in a distributed manner. Similarly, $P_{kj}=1$ indicates that the kthk-th sequence is assigned to device $j$, implying $D_{k}=0$. Given a micro-batch comprising $K$ sequences with lengths $S_{k}$ ($k=(0,\ldots,K-1)$), BucketSize $C$ and CP degree $N$, the scheduling process of DACP can be formulated as follows:
@@
$\text{Time}_{j}=\max\left(T_{comm}\left(\text{V}\right),T_{comp}\left(\text{Local}_{j}\right)\right)+T_{comp}\left(\text{Dist}\right),\qquad\forall j$ (2)
$\text{Local}_{j}=\sum_{k}\text{FLOPs}\left(P_{kj}\cdot S_{k}\right),\qquad\forall j$ (3)
@@
$\sum_{j}P_{kj}+D_{k}=1,\qquad\forall k$ (6)
$\sum_{k}S_{k}\cdot P_{kj}+\frac{D_{k}\cdot S_{k}}{\text{N}}\leq C,\qquad\forall j$ (7)
Here, our optimization goal is to find the optimal $D$ and $P$ to minimize $TDACP$, which represents the total duration in one micro-batch. As shown in Equation 1, $TDACP$ is determined by the maximum execution time $\text{Time}_{j}$ across all CP ranks $j$. Specifically, as described in Equation 2, $\text{Time}_{j}$ consists of two components: (1) the overlapping term, defined as the maximum of the communication time $T_{comm}(\text{V})$ and the computation time $T_{comp}(\text{Local}_{j})$ for local sequences, and (2) the computation time $T_{comp}(\text{Dist})$ for distributed sequences. Here, $T_{comm}$ depends on the communication volume V, as modeled in Equation 5. Similarly, $T_{comp}$ utilizes the results from Equations 3 and 4, which compute the FLOPs for local sequences on CP rank $j$ and distributed sequences, respectively. Finally, Equation 6 ensures the completeness of data scheduling, while Equation 7 enforces the memory constraint.
@@ 4.2 Global Data Scheduling
First, the heterogeneous sequence length distribution also leads the to severe load imbalance across different micro-batches, resulting in the sub-optimal training efficiency in Long-SFT scenarios. Second, to achieve maximum performance gains in DACP, meticulous micro-batching strategy is essential. For example, pairing long and short sequences with appropriate memory pressure can expand the valid scheduling space for DACP. Specifically, micro-batches with large total sequence lengths increase the risk of OOMs and limit the optimizations in DACP, such as selective sharding. In contrast, micro-batches with small total sequence lengths introduce GPU under-utilization, degrading the end-to-end performance. Therefore, as shown in Figure 2(b), Skrull employs Global Data Scheduling (GDS), which derives the optimal micro-batching strategy from the global batch. We limit the scheduling scope to the global batch because it represents the maximum scope that maintains mathematical equivalence for mainstream optimizers such as Adam [optimizer_adam_2014] and AdamW [optimizer_adamw_2017].
@@
$\sum_{ij}B_{kij}=1,\qquad\forall k$ (9)
$\sum_{ki}B_{kij}*S_{k}\leq C*N,\qquad\forall j$ (10)
$\text{Time}_{ij}=\textbf{TDACP}\left(B_{kij}*S_{k}\right),\qquad\forall k$ (11)
@@ 4.3 Efficient Online Scheduling
Although some solvers like [SCIP_MILP_Sovler] can derive the optimal scheduling plan, its long solving time makes it impractical for scheduling during runtime. To achieve online scheduling during Long-SFT, we resort to design lightweight heuristic scheduling algorithm. Notably, our scheduling algorithm is integrated into the DataLoader and introduces near-zero overhead to the training process.
@@
Unfortunately, we cannot balance the computation and memory at the same time. The reason is that, after applying FlashAttention [flashattention_v1_2022, flashattention_v2_2024], the correlation between computation complexity and and sequence length ($n$) is $O(n^{2})$, however, the correlation between memory is $O(n)$. Moreover, with the sequence length increasing, the portion of Attention module gradually dominates the computation load, making it more difficult to balance the computation and memory. Worse still, the model configuration (e.g., KV heads, hidden size) also impacts. Due to the limited page, we list the details in Appendix C.
@@ 4.3.2 Heuristics
We first summarize three principles of algorithm design in DACP. (i) Avoid sharding: We strive to avoid sequence sharding and assume that all sequences will be handled locally first. (ii) Prioritize computation: We prioritize balancing computation over memory to achieve better performance. (iii) Roll-back mechanism: We continuously monitor the estimated memory consumption and revert decisions when necessary. The roll-back mechanism guarantees the memory constrains outlined in Equation 7 and Equation 10, while enabling more aggressive scheduling attempts based on (i) and (ii). Our heuristic for DACP is listed in Algorithm 1. Given a micro-batch containing $K$ sequences with lengths $S[K]$ and a predefined BucketSize $C$, the algorithm outputs the sequence classification and assignment results in the form of an array retret. In this array, a value of -1 at the i-th position indicates that the i-th sequence is to be sharded, while a value $v=(0,\ldots,ws-1)$ indicates that the i-th sequence is assigned to CP rank $v$ entirely. To better balance computation while ensuring memory constrains, we maintain two arrays during DACP scheduling: RemainBucket $RB$ and Loads $L$, which represent the current memory budget and computation load, respectively. We first sort the sequences in ascending order. For the each sequence, we sequentially assign it to the bucket (as well as CP rank) with minimum $L$ to avoid sharding and prioritize balancing computation (line 6-8). If the bucket cannot accommodate the sequence, we attempt to assign it to the bucket with the maximum $RB$ to avoid sharding (line 10-12). If both attempts fail, we classify the sequence as a distributed sequence and attempt to shard it (line 14-16). However, if the bucket with minimum of $RB$ cannot handle the sub-sequence after sharding, this indicates that the earlier process incorrectly classified inappropriate sequences as local sequences within this bucket. To address this, we employ a roll-back mechanism (line 18 and Appendix B.1). This mechanism identifies a local sequence in the bucket, shards it to reduce memory pressure, and resumes the assignment process. If the roll-back fails due to the absence of local sequences in the bucket, we return a DACP scheduling error. In such cases, GDS will also revert the batching plan (see the Section 4.3.2). Notably, every assignment updates $RB$ and $L$ through the predefined functions UpdateLocal and UpdateAll. The details of these functions including RollBack are further elaborated in Appendix B.1.
@@
1:SeqNum $K$, SeqLens $S[K]$, BucketSize $C$, CP degree NN
@@
4:for $i=0$ to N1N-1 do
5:  $RB[i]{\leftarrow}C,L[i]{\leftarrow}0$ $\triangleright$ Initialization
@@
7:for $i=0$ to K1K-1 do
8:  $t\leftarrow\text{argmin}(L)$ $\triangleright$ Find rank t with minimum workload
@@
18:      $ret[i]{\leftarrow}{-}1$, UpdateAll(i)(i) $\triangleright$ Distribute the sequence
@@
20:      Assert RollBack(t,RB,L)(t,RB,L)
21:      $i{\leftarrow}i{-}1$ $\triangleright$ Roll-back to avoid OOMs
@@
Algorithm 2 demonstrates the heuristic scheduling algorithm of GDS. Given a global batch containing $K$ sequences with lengths $S[K]$, DP world size $ws$ and DP rank dp_rankdp\_rank, the algorithm returns the scheduling result mbsmbs, which consists of multiple micro-batches as inputs for Algorithm 1. We summarize three principles in our algorithm design. (i) Prioritize computation: We prioritize balancing computation across DP workers. To achieve this, we estimate the FLOPs (Appendix C.2) and employ a bin-packing algorithm to balance computational workloads at a coarse granularity (line 1). (ii) Pair long and short sequences: We sort the sequences within each DP rank and batch them in an interleaved manner (line 7). This approach ensures that long sequences are assigned more evenly across micro-batches. Additionally, each micro-batch contains several short sequences, enhancing both task overlapping and load balancing. (iii) Improve memory utilization: We estimate the total memory requirements and try to improve the concurrency with less number of micro-batches. Thanks to the roll-back mechanism (line 8), this method maximizes memory utilization while not increase the risk of OOMs. As shown in line 5, we gradually increase the number of micro-batches if the scheduling fails and requires a roll-back.
@@
1:SeqNum $K$, SeqLens $S[K]$, BucketSize $C$, CP degree $N$, DP WorldSize wsws, DP_Rank dp_rankdp\_rank
2:Micro-batches mbsmbs
[…]
5:Sort(SubsetSubset, ascending=True)
@@
8:  for $j\leftarrow 0$ to initinit do
9:    $mbs.\text{append}(Subset[j::init])$ $\triangleright$ Pair long and short sequences